OpenStudy (anonymous):

Choose the correct graph of the linear system. Then enter the solution of the system. x-y=1 x+y=-5

OpenStudy (uri):

Add them x-y=1 x+y=-5 ===== 2x=-4 x=-4/2 x=-2

OpenStudy (anonymous):

How did you get the 2x?

OpenStudy (anonymous):

OK That's how I solve for X To solve for Y do I subtract 1 from -5 or -5 from 1 or neither?

OpenStudy (uri):

Nah to solve for y put the value of x in eq no.2 x+y=-5 -2+y=-5 y=-5+2 y=-3

OpenStudy (anonymous):

Thanks! I forgot that to solve for Y you have to solve for X first, one more question. Where did you get the 2x from? is it because you added x-y=1 and x+y=-5+ Trying to get these liner equations through my head.

OpenStudy (uri):

x+x =2x

OpenStudy (uri):

yes and -y +y will cancelled.

OpenStudy (anonymous):

Thanks again!

OpenStudy (uri):

yw :) someone who wants to learn.

OpenStudy (anonymous):

@uri Closed the question too soon! I figured out the graph (the third option) but Now I have to answer What is the solution of the linear system? What do I put there?

OpenStudy (uri):

Put where? c:

OpenStudy (anonymous):

In the question "What is the solution of the linear system?" c:

OpenStudy (anonymous):

*cool smiley

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

And thank you for all your help!

OpenStudy (uri):

(-2,-3)

OpenStudy (anonymous):

OH OK! I feel dumb :P